OpenStudy (anonymous):

what is the quiotient of 5/6 and 1/2

OpenStudy (aakashsudhakar):

Dividing by a fraction is the same by multiplying by its reciprocal. Therefore... (5/6)/(1/2) = (5/6) * 2 = 10/6 = 5/3 The answer is five-thirds, or 5/3, or roughly 1.6666666667.